# Large Language Models Synergize with Automated Machine Learning

### v1

@@ Abstract
Recently, code generation driven by large language models (LLMs) has become increasingly popular. However, automatically generating code for machine learning (ML) tasks still poses significant challenges. This paper explores the limits of program synthesis for ML by combining LLMs and automated machine learning (autoML). Specifically, our goal is to fully automate the code generation process for the entire ML workflow, from data preparation to modeling and post-processing, utilizing only textual descriptions of the ML tasks. To manage the length and diversity of ML programs, we propose to break each ML program into smaller, manageable parts. Each part is generated separately by the LLM, with careful consideration of their compatibilities. To implement the approach, we design a testing technique for ML programs. Furthermore, our approach enables integration with autoML. In our approach, autoML serves to numerically assess and optimize the ML programs generated by LLMs. LLMs, in turn, help to bridge the gap between theoretical, algorithm-centered autoML and practical autoML applications. This mutual enhancement underscores the synergy between LLMs and autoML in program synthesis for ML. In experiments across various ML tasks, our method outperforms existing methods in 10 out of 12 tasks for generating ML programs. In addition, autoML significantly improves the performance of the generated ML programs. In the experiments, our method, Text-to-ML, achieves fully automated synthesis of the entire ML pipeline based solely on textual descriptions of the ML tasks.
@@ 1.1 Background, challenges, and motivation
Program synthesis is the automatic generation of computer programs from high-level specifications Manna & Waldinger (1971). In recent years, autoregressive large language models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Anil et al. (2023) have demonstrated impressive performance in program synthesis tasks Austin et al. (2021); Li et al. (2022); Shinn et al. (2023). However, existing LLM-based program synthesis methods primarily focus on traditional coding problems, such as interview questions for software engineering positions Austin et al. (2021); Chen et al. (2023b); Madaan et al. (2023), while the generation of machine learning (ML) programs remains largely unexplored (Figure 2). The challenges associated with synthesizing ML programs stem from their length, diversity, and complexity in testing phase:
@@
  • Length: The program for a complete ML pipeline for a given task consists of several task-specific components: (1) data preparation, such as data loading, cleaning, encoding, and feature engineering; (2) modeling, which implements the ML algorithm; and (3) post-processing, such as performance evaluation and decoding the results (Figure 1). These components collectively demand a lengthy output from LLMs. For instance, the validly generated code for the ML task depicted in Figure 1 consists of an average of $114\pm 32$ lines, in stark contrast to the average $6.3\pm 5.0$ lines of the reference solutions in the HumanEval benchmark, which addresses traditional coding problems Chen et al. (2021). For LLMs, the length of the output sequence strongly correlates with the difficulty of the generation process Anil et al. (2022); Newman et al. (2020).
@@
  • Diversity: To achieve the automatic generation of the program for the entire pipeline, many task-specific components of the ML program should be generated during the process, rather than relying on an inventory of pre-written modules Surís et al. (2023); Gupta & Kembhavi (2023); Lu et al. (2023). Depending on pre-written modules for a wide range of tasks is often impractical (Hollmann et al., 2023), especially given the vast combinations of requirements characteristic across different ML tasks, such as the structure of dataset files, dataset features, and desired output format.
@@
  • Complexity in testing phase: Evaluation of programs for traditional coding problems often relies on manually written white-box tests (Chen et al. (2023b); Li et al. (2022)), such as unit tests Myers et al. (2004). Unit tests verify the correctness of individual code components, such as functions or classes, by comparing their actual behavior with the expected behavior based on specific inputs and outputs. For ML programs, unit units are insufficient and costly to create Tian et al. (2018); Wang et al. (2021). From the perspective of both software testing and ML, syntactically valid ML programs should be further examined based on numerical performance metrics, such as Top-1 accuracy or mean absolute error, and a consistent evaluation procedure for semantic accuracy Kim et al. (2019); Braiek & Khomh (2020); Saha et al. (2022); Cambronero & Rinard (2019).
ML heavily impacts modern society, with applications spanning healthcare, finance, manufacturing, and beyond. Implementing ML typically demands substantial expertise and effort, which limits ML’s accessibility for many potential users. While many works address automating ML, the full automation of ML tasks at the code level remains an unresolved challenge (see Sections 2.2, 2.3, and A.2). In this paper, we explore the potential of program synthesis for ML. During this exploration, we aim to automate the full ML pipeline at the code level, reducing the human workload to mere textual descriptions of the tasks. Our study focuses on automating supervised ML tasks, encompassing traditional algorithms such as random forests and SVM, as well as deep learning domains including computer vision (CV) and natural language processing (NLP).
@@
Figure 3: Method overview. Initially, based on the task description, the LLM generates the modules. The modules are verified by unit tests to ensure inter-compatibility. With the compatibility, they are automatically integrated to form an executable program. This process iterates to produce multiple program candidates. Finally, autoML selects the best program and hyperparameters based on the evaluation metrics.
1.2 Sequence generation framework for autoregressive LLMs
@@
Recently, many studies have leveraged autoregressive LLMs for sequence generation tasks. For example, Chain-of-Thought Wei et al. (2022) enhances reasoning performance by incorporating input-output examples within the input sequence, while Reflexion Shinn et al. (2023) utilizes a reinforcement learning strategy to boost generation performance by reflecting on feedback. Despite these advancements, these frameworks still struggle with generating ML programs (see Section 6). To manage the length and diversity in ML programs, we introduce a novel sequence generation framework for autoregressive LLMs, Contextual Modular Generation (Algorithm 1). As a form of compositional reasoning Andreas et al. (2016), this framework breaks down the programs into multiple smaller modules, with each module being generated in a separate process and representing a distinct code component of the overall program (Figures  1 and 3).
@@
Unlike traditional compositional approaches Andreas et al. (2016); Surís et al. (2023); Gupta & Kembhavi (2023); Lu et al. (2023), Contextual Modular Generation is distinctive for two features: (1) instead of hinging on a fixed inventory of pre-written modules (e.g., selecting and reusing code from a repository of pre-written codes for available feature engineering methods), the framework dynamically generates the essential modules tailored to each specific task. In our study, the generated modules include data preparation, modeling, and post processing. (2) the generated modules are readily compatible without requiring additional sequence generation steps from LLMs for combination. This feature contributes to scalability by ensuring linear growth in generation complexity over program length (Theorem 3.8).
1.3 Testing technique for ML programs
@@
1.4 Incorporation with autoML
Figure 4: AutoML in our method. First, the LLM generates the code for multiple pipelines as candidate solutions. Next, the optimization algorithm chooses a solution. This chosen solution is then trained and evaluated, providing feedback to improve selections in subsequent iterations.
AutoML aims to automatically select the optimal pipeline and hyperparameters for ML tasks Thornton et al. (2013). For an ML task, an AutoML process begins by designating a solution and then proceeds to data processing and model training based on the solution. Subsequently, the solution is evaluated, providing feedback for future iterations to designate more promising solutions (Figure 4). Typically, AutoML utilizes optimization algorithms, such as evolutionary computation or reinforcement learning, to designate the solutions. In our study, the decision variables in the optimization (i.e., the elements of the solution) include choices and hyperparameters regarding data preparation methods, ML models, and post-processing methods implemented by the ML program. The objective function (i.e., the goal of the solution) is the prediction performance of the ML program, measured by the numerical metrics.
Due to the aforementioned complexity involved in testing ML programs, existing works in program synthesis for ML often employ autoML methods (see Section 2.3). Our autoML method adapts the strategies in Öztürk et al. (2022) by incorporating zero-cost (ZC) proxies Mellor et al. (2021) (see Section C.5). These proxies predict the performance of the trained neural networks using at most a single forward/backward propagation pass, which significantly reduces the cost from evaluating (training and testing) the models.
@@ 1.5 Summary
From a sequence generation perspective, our work explores the resolution of ML problems as a multi-step sequence to sequence (seq2seq) task Sutskever et al. (2014). The input sequence is the textual task description. The output sequence is the corresponding optimized program for the entire pipeline. Intermediary steps include sub-seq2seq tasks based on LLMs and optimization via autoML. Our method, Text-to-ML, seamlessly integrates all intermediary steps into a single fully autonomous process (Figure 3 and Algorithm 2). Our contributions are summarized as follows:
  • We explore the limits of program synthesis for ML. Qualitatively, for 12 supervised ML tasks over traditional data, CV, and NLP, Text-to-ML generates the full optimized pipeline in a multi-step seq2seq process. In the exploration, we offer a new perspective for combining LLMs and autoML.
  • We propose a sequence generation framework Contextual Modular Generation for autoregressive LLMs, with a formal analysis of scalability over sequence length (Theorem 3.5 and Theorem 3.8). To implement the framework for the generation of ML programs, we devise a novel testing technique based on progenitor testing protocols (see Section 4) and LLMs.
  • For 10 out of 12 ML tasks across modalities, Contextual Modular Generation outperforms existing methods. Furthermore, autoML substantially improves the performance of the generated programs. In addition, we find ZC proxies speed up the search with both pretrained CV and NLP models.
@@ 2.1 Large language models
Compositional reasoning decomposes a reasoning process into multiple intermediate steps Andreas et al. (2016). Compared to traditional compositional reasoning Andreas et al. (2016); Surís et al. (2023); Gupta & Kembhavi (2023); Lu et al. (2023), our method generates modules during the reasoning process. In this aspect, a similar method is Le et al. (2023). Both methods decompose the generation of the output sequence into multiple generations of smaller sequences: $Y^{1},Y^{2},\ldots,Y^{J}$, where each $Y^{j}=(y^{j}_{1},y^{j}_{2},\ldots,y^{j}_{k})$. For each $Y^{j}$, let the correpsonding input sequence be $X^{j}$, then $p({Y^{j}}|X^{j})=\prod_{n=1}^{N}p_{\lambda}(y^{j}|y^{j}_{<n},X^{j})$.
A major challenge for generating the modules is whether $Y^{1},$ $Y^{2},\ldots,Y^{J}$ form a valid solution. Le et al. (2023) addresses the issue by incorporating the modules as the input sequence in an additional generation step to adapt the modules for the final output. However, Le et al. (2023) mainly targets traditional coding problems. For ML programs, this additional generation step is less effective (Table 1), due to the length of the programs. By comparison, our method requires no extra generation step for combining the modules.
@@ 2.2 Automated machine learning
AutoML methods typically involve an optimization algorithm and a search space (i.e., a set of candidate solutions) Zoph & Le (2016); Liu et al. (2018). Öztürk et al. (2022) explores the search space of pretrained deep learning models. Our autoML method adapts the strategies in Öztürk et al. (2022). In addition, we incorporate ZC proxies from Mellor et al. (2021); Tanaka et al. (2020) in our method, which accelerates the optimization for deep learning tasks by predicting the performance of the models.
Very recently, several works have explored the application of LLMs for autoML, however, none of the works achieve automatic generation of the entire pipeline. For a qualitative comparison, see Appendix A.2.
@@ 2.3 Software engineering
Program synthesis is a longstanding objective within the arctificial intelligence research community Manna & Waldinger (1971). Recently, there has been a surge of interest in ML for program synthesis, particularly with LLMs Austin et al. (2021); Li et al. (2022). Conversely, program synthesis for ML is also an emerging topic and often involves autoML. Real et al. (2020) evolves programs for ML models. Saha et al. (2022); Cambronero & Rinard (2019) construct the programs for the modeling and part of data preparation for tasks about tabular data. Leveraging LLMs, our research expands the scope of program synthesis for ML by targeting automatic generation of the entire pipeline across task modalities.
Software testing techniques Myers et al. (2004) can assist program synthesis in evaluating the programs. Software testing for ML programs is a challenging topic Braiek & Khomh (2020). Pei et al. (2017) and Tian et al. (2018) address the challenge for computer vision tasks. Recently, Chen et al. (2022); Shinn et al. (2023) employ LLMs to automatically generate tests for traditional coding problems. Based on LLMs, the testing technique in our study is designed for implementing the Contextual Modular Generation framework. Compared to existing works, our technique is uniquely applicable for testing ML programs at a modular level and across task modalities (traditional algorithms, CV, and NLP).

### v2

@@ Abstract
Recently, program synthesis driven by large language models (LLMs) has become increasingly popular. However, program synthesis for machine learning (ML) tasks still poses significant challenges. This paper explores a novel form of program synthesis, targeting ML programs, by combining LLMs and automated machine learning (autoML). Specifically, our goal is to fully automate the generation and optimization of the code of the entire ML workflow, from data preparation to modeling and post-processing, utilizing only textual descriptions of the ML tasks. To manage the length and diversity of ML programs, we propose to break each ML program into smaller, manageable parts. Each part is generated separately by the LLM, with careful consideration of their compatibilities. To ensure compatibilities, we design a testing technique for ML programs. Unlike traditional program synthesis, which typically relies on binary evaluations (i.e., correct or incorrect), evaluating ML programs necessitates more than just binary judgments. Therefore, we further assess ML programs numerically and select the optimal programs from a range of candidates using AutoML methods. In experiments across various ML tasks, our method outperforms existing methods in 10 out of 12 tasks for generating ML programs. In addition, autoML significantly improves the performance of the generated ML programs. In experiments, given the textual task description, our method, Text-to-ML, generates the complete and optimized ML program in a fully autonomous process.
@@ 1.1 Background, challenges, and motivation
Program synthesis is the automatic creation of computer programs that satisfy high-level specifications (Manna & Waldinger, 1971). In recent years, autoregressive large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Anil et al., 2023) have demonstrated impressive performance in program synthesis tasks (Austin et al., 2021; Li et al., 2022; Shinn et al., 2023). However, existing LLM-based program synthesis methods primarily focus on traditional coding problems, such as interview questions for software engineering positions (Austin et al., 2021; Chen et al., 2023b; Madaan et al., 2023), while the synthesis of machine learning (ML) programs remains largely unexplored (Figure 2). The challenges associated with program synthesis for ML stem from their length, diversity, and complexity in testing phase:
@@
  • Length: The first step in our study is to generate an entire ML program, which consists of three task-specific components: (1) data preparation, such as data loading, cleaning, encoding, and feature engineering; (2) modeling, which implements an ML model, such as BERT (Devlin et al., 2018); and (3) post-processing, such as performance evaluation and decoding the results (Figure 1). These components collectively demand a lengthy output from LLMs. For instance, the validly generated code for the ML task depicted in Figure 1 consists of an average of $114\pm 32$ lines, in stark contrast to the average $6.3\pm 5.0$ lines of the reference solutions in the HumanEval benchmark, which addresses traditional coding problems (Chen et al., 2021). For LLMs, the output length strongly correlates with the difficulty of the generation process (Anil et al., 2022; Newman et al., 2020).
@@
  • Diversity: To achieve the automatic generation of the program for the entire ML program, many task-specific components of the ML program should be generated during the process, rather than relying on an inventory of pre-written modules (Surís et al., 2023; Gupta & Kembhavi, 2023; Lu et al., 2023). Depending on pre-written modules for a wide range of tasks is often impractical (Hollmann et al., 2023), especially given the vast combinations of requirements characteristic across different ML tasks, such as the structure of dataset files, dataset features, and desired output format.
@@
  • Complexity in testing phase: Evaluation of programs for traditional coding problems often relies on manually written white-box tests (Chen et al., 2023b; Li et al., 2022), such as unit tests (Myers et al., 2004). Unit tests verify the correctness of individual code components, such as functions or classes, by comparing their actual behavior with the expected behavior based on specific inputs and outputs. For ML programs, unit tests are insufficient and costly to create (Tian et al., 2018; Wang et al., 2021b). From the perspective of both software testing and ML, syntactically valid ML programs should be further examined based on numerical performance metrics, such as Top-1 accuracy or mean absolute error, and a consistent evaluation procedure for semantic accuracy (Kim et al., 2019; Braiek & Khomh, 2020; Saha et al., 2022; Cambronero & Rinard, 2019).
Notably, the complexity in testing phase entails a novel form of program synthesis, setting it apart from traditional program synthesis. In traditional contexts, program synthesis often concludes with code generation and perhaps testing through binary evaluation procedures, such as unit tests. However, ML programs necessitate numerical evaluations, extending the synthesis process beyond just code generation (Kim et al., 2019; Braiek & Khomh, 2020; Saha et al., 2022; Cambronero & Rinard, 2019).
Firstly, how to ensure that each generated program is compatible with the code of a consistent and fair training, testing, and evaluation procedure? Secondly, how to effectively utilize numerical feedback involving performance scores? A numerical optimization algorithm efficiently addresses these challenges. Consequently, for ML programs, program synthesis often necessitates optimization. Real et al. (2020); Saha et al. (2022); Cambronero & Rinard (2019) also explore this form of program synthesis (see Section 2.3). Similar to these studies, we also utilize automated machine learning (autoML) as the numerical optimization step.
@@
Figure 3: Text-to-ML: a three-phase process of generation, testing, and optimization. Initially, with the task description, the LLM generates the modules. The modules are verified by unit tests to ensure inter-compatibility. With the compatibility, they are automatically integrated. Then, iterate to produce multiple candidate programs. Finally, autoML selects the best program and associated hyperparameters.
ML heavily impacts modern society, with applications spanning healthcare, finance, manufacturing, and beyond. Implementing ML typically demands substantial expertise and effort, which limits ML’s accessibility for many potential users. Many studies work on automating ML utilizing LLMs or autoML, however, the full automation of ML tasks at the code level remains an unresolved challenge (see Sections 2.2 and 2.3 and Appendix A.2). In this paper, we explore the limits of program synthesis for ML by combining LLMs and autoML. During this exploration, we aim to automate the generation and optimization of the entire ML program, reducing the human workload to mere textual descriptions of the tasks. Our study focuses on synthesizing the programs involving supervised ML tasks, encompassing traditional algorithms such as random forests and SVM, as well as deep learning domains including computer vision (CV) and natural language processing (NLP).
1.2 Automatic code generation
@@
Recently, many studies have leveraged autoregressive LLMs for sequence generation tasks. For example, Chain-of-Thought (Wei et al., 2022) enhances reasoning performance by incorporating input-output examples within the input sequence, while Reflexion (Shinn et al., 2023) utilizes a reinforcement learning strategy to boost generation performance by reflecting on feedback. Despite these advancements, these frameworks still struggle with generating ML programs (see Table 1). To manage the length and diversity in ML programs, we introduce a novel sequence generation framework for autoregressive LLMs, Contextual Modular Generation (Algorithm 1). As a form of compositional reasoning (Andreas et al., 2016), this framework breaks down the programs into multiple smaller modules, with each module being generated in a separate process and representing a distinct code component of the overall program (Figures  1 and 3).
@@
Unlike traditional compositional approaches (Andreas et al., 2016; Surís et al., 2023; Gupta & Kembhavi, 2023; Lu et al., 2023), Contextual Modular Generation is distinctive for two features: (1) instead of hinging on a fixed inventory of pre-written modules (e.g., selecting and reusing code from a repository of pre-written codes for available feature engineering methods), the framework dynamically generates the essential modules tailored to each specific task. In our study, the generated modules include data preparation, modeling, and post processing. (2) the generated modules are readily compatible without requiring additional sequence generation steps from LLMs for combination. This feature contributes to scalability by ensuring linear growth in generation complexity over program length (Theorem 3.8).
1.3 Automatic unit testing
@@
1.4 Automatic code optimization
Figure 4: The autoML step in our method. Each solution is a program that implements a candidate combination of data preparation, modeling, and post-processing algorithms and associated hyperparameters. First, the optimization algorithm chooses a solution. This chosen solution is then trained and evaluated, which provides feedback to improve selections in subsequent iterations.
The last step of our method is autoML (Figure 3). AutoML is the process of automatically selecting algorithms and hyperparameters for ML tasks (Thornton et al., 2013). For a task, an AutoML process begins by designating a solution and then proceeds to data processing and model training based on the solution. Subsequently, the solution is evaluated, providing feedback for future iterations to designate more promising solutions (Figure 4). Typically, AutoML utilizes optimization algorithms, such as evolutionary computation or reinforcement learning, to designate the solutions. In our study, the decision variables in the optimization (i.e., the elements of the solution) include choices and hyperparameters regarding data preparation methods, ML models, and post-processing methods implemented by the ML program. The objective function (i.e., the goal of the solution) is the prediction performance of the ML program, measured by the numerical metrics.
Due to the aforementioned complexity involved in testing ML programs, existing works in program synthesis for ML often employ autoML methods (see Section 2.3). The autoML step in our method adapts the strategies in Öztürk et al. (2022) by incorporating zero-cost (ZC) proxies (Mellor et al., 2021) (see Appendix C.5). These proxies predict the performance of the trained neural networks using at most a single forward/backward propagation pass, which significantly reduces the cost from evaluating (training and testing) the models.
@@ 1.5 Summary
Figure 5: Program synthesis for ML. (1) black rectangle: the goal of this study. (2) red rectangles: challenges. (3) green rectangles: methods or features of methods. (4) solid rectangles: novel goals, methods, or features introduced in our study. (5) dashed rectangles: existing challenges or methods.
Figure 5 illustrates the relation among the goals, challenges, and methods in our study. Our contributions are summarized as follows:
  • We explore the limits of program synthesis for ML. Qualitatively, for 12 supervised ML tasks over traditional data, CV, and NLP, Text-to-ML generates the code of the end-to-end, optimized program in a fully automatic process, given textual descriptions of the tasks (Figure 1 and Section 6.2). In the exploration, we offer a new perspective for combining LLMs and autoML. We identify a synergy in this combination that enhances program synthesis for ML (see Section 7.1).
  • We propose a sequence generation framework Contextual Modular Generation for autoregressive LLMs, with a theoretical analysis of scalability over sequence length (Theorem 3.5 and Theorem 3.8). To ensure compatibility among the modules, as demanded by the framework, we devise a novel testing technique based on progenitor testing protocols (see Section 4) and LLMs.
  • For 10 out of 12 ML tasks across modalities, Contextual Modular Generation outperforms existing methods (Section 6.2). Furthermore, autoML substantially improves the performance of the generated programs (Section 6.4). In addition, we find ZC proxies speed up the search with both pretrained CV and NLP models (Section 6.4).
@@ 2.1 Large language models
Compositional reasoning decomposes a reasoning process into multiple intermediate steps (Andreas et al., 2016). Compared to traditional compositional reasoning (Andreas et al., 2016; Surís et al., 2023; Gupta & Kembhavi, 2023; Lu et al., 2023), our method generates modules during the reasoning process. In this aspect, a similar method is Le et al. (2023). Both methods decompose the generation of the output sequence into multiple generations of smaller modules: $Y^{1},Y^{2},\ldots,Y^{J}$, where each $Y^{j}=(y^{j}_{1},y^{j}_{2},\ldots,y^{j}_{k})$. For each $Y^{j}$, let the correpsonding input sequence be $X^{j}$, then $p({Y^{j}}|X^{j})=\prod_{n=1}^{N}p_{\lambda}(y^{j}|y^{j}_{<n},X^{j})$.
A major challenge for generating the modules is whether $Y^{1},$ $Y^{2},\ldots,Y^{J}$ form a valid solution. Le et al. (2023) addresses the issue by incorporating the modules as the input sequence in an additional generation step to adapt the modules for the final output. However, Le et al. (2023) mainly targets traditional coding problems. For ML programs, this additional generation step is less effective (Table 1), due to the length of the programs. By comparison, our method automatically ensures the compatibility of generated modules and requires no extra generation step for combination.
@@ 2.2 Automated machine learning
AutoML methods typically involve an optimization algorithm and a search space (i.e., a set of candidate solutions) (Zoph & Le, 2016; Liu et al., 2018). Öztürk et al. (2022) explores the search space of pretrained deep learning models. Our autoML method adapts the strategies in Öztürk et al. (2022). In addition, we incorporate ZC proxies from Mellor et al. (2021) and Tanaka et al. (2020) in our method, which accelerates the optimization for deep learning tasks by predicting the performance of the models.
Very recently, several works have explored the application of LLMs for autoML, however, none of the works achieve automatic generation of the entire ML program. For a qualitative comparison, see Appendix A.2.
@@ 2.3 Software engineering
Program synthesis is a longstanding objective within the arctificial intelligence research community (Manna & Waldinger, 1971). Recently, there has been a surge of interest in ML for program synthesis, particularly with LLMs (Austin et al., 2021; Li et al., 2022). Conversely, program synthesis for ML is also an emerging topic and often involves autoML. Real et al. (2020) evolves programs for ML models. Saha et al. (2022); Cambronero & Rinard (2019) construct the programs for the modeling and part of data preparation for tasks about tabular data. Leveraging LLMs, our research expands the scope of program synthesis for ML by targeting the entire ML program across task modalities.
Software testing techniques (Myers et al., 2004) can assist program synthesis in evaluating the programs. Software testing for ML programs is a challenging topic (Braiek & Khomh, 2020). Pei et al. (2017) and Tian et al. (2018) address the challenge for computer vision tasks. Recently, Chen et al. (2022); Shinn et al. (2023) employ LLMs to automatically generate tests for traditional coding problems. Based on LLMs, the testing technique in our study is designed for implementing the Contextual Modular Generation framework. Compared to existing works, our technique is uniquely applicable for testing ML programs at a modular level and across task modalities (traditional algorithms, CV, and NLP).